OpenStudy (anonymous):

Which of the following is a solution of x2 - 7x = -5?

OpenStudy (anonymous):

A. \[\frac{ 7\pm \sqrt{69} }{ 2 }\] B. \[\frac{ 7\pm \sqrt{29} }{ 2 }\] C. \[\frac{ -7\pm \sqrt{29} }{ 2 }\] D. \[\frac{ -7\pm \sqrt{69} }{ 2 }\]

OpenStudy (welshfella):

a =1 , b = -7 and c = 5

OpenStudy (anonymous):

Ok. I'm gonna see if I can do it first.

OpenStudy (anonymous):

I know its either B or C. I'm pretty ure it's B tho

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

its -(-7) = 7 so yes its B

OpenStudy (anonymous):

Ok thank you again :)

OpenStudy (welshfella):

yw